OpenStudy (anonymous):

Please Help, Ill give medals & Fans Q1. Solve the following equation. | 2v + 5 | = 9 A.2 or -7 B.-2 and 7 C.2 D.-7 Q2. Solve the following equation. | 5b | - 7 = 13 A.4 B.5 C.2 D.3

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (anonymous):

that site should help, plus if you sign up it will show you how to solve the problem

OpenStudy (imstuck):

The first one comes to 2v+ 5 =9 and the second solution for that is 2v + 5 = -9. Solve both of those for v.

OpenStudy (anonymous):

Question 1 is c. :)

OpenStudy (anonymous):

Thanks Guys!! (: @christena710 @IMStuck @William.deems

OpenStudy (anonymous):

no problem ;)

OpenStudy (lakota1999):

one comes to 2v+ 5 =9 and second onethat is 2v + 5 = -9. Solve both of those for v.